What is the change in internal energy (in J) of a system that releases 675 J of thermal energy to its surroundings and has 9.90 × 102 cal of work done on it? Give your answer in scientific notation.

 



 

1
Expert's answer
2020-10-23T09:15:16-0400

If system releases heat q is negative,

If work done on the system w is positive

9.90 × 102 cal x 4.184 J = 4142.16 J

Δ E = q + w = -675 J + 4142.16 J = 3467.16 J = 3.4 x 103 J